# A Quarter of a Century of Neuromorphic Architectures on FPGAs - an Overview

### v2

@@ 3.4. Class 3 (Traits: asynchronous network update)
Early implementations. Hellmich et al. (2004,2005) (Hellmich and Klar, 2004; Hellmich et al., 2005) presented SNN Emulation Engine (SEE), which they claimed it could support up to 50 million IF neurons and 800 million synapses. The architecture organization relied on three FPGAs responsible for: (i)network configuration, monitoring and administration, (ii) computation of the target neurons addresses and (iii) updating neuron potentials and weight updates. A set of three PEs was implemented on the third device connected to three bespoke SDRAMs. The neuron update used a Bulirsch-Stoer integration method.
Figure 6. Simplified diagram of a Class 3 NMA.
@@
Ros et al. (2006) (Ros et al., 2006) presented RT-Spike, which was a SW/HW hybrid architecture for realizing SNNs with SRM-model-based neurons. The authors chose a 14b fixed-point number representation. They tested the system with a network of 1024 neurons on four PEs and achieved real-time operation on Xilinx Virtex-2000E at 25 MHz, with 4.16 million updates of the network state per second. Glackin et al. (2009) (Glackin et al., 2009) presented Reconfigurable Architecture, which aimed at modeling the biological functions of the brain. The architecture comprised a set of PEs simultaneously simulating up to 250 neuron with COBA LIF and SRM neuron models support. Reconfiguring the network was performed on-the-fly, by changing the values in the BRAM memory via a custom DMA core. The implemented system performed edge detection on 512x512 px images on a Xilinx Virtex-4 device with four PEs, using an SCNN with 1.05 million neurons and 52.4 million synapses, reaching 1.05 seconds of inference time per image.
@@
Early modern implementations. Yang et al. (2011) (Yang et al., 2011) presented an NMA for explaining the computational capability of the early stage of the primate visual system. The input for the topology was a gray image transformed into a temporal firing rate in a given period. The architecture used four FPGAs, one per every image orientation feature, with 64 TDM PEs per device, allowing for one million LIF neurons and two million synapses. It was implemented on the Xilinx Virtex-4 device and clocked at 200 MHz. Neil et al. (2014) (Neil and Liu, 2014) presented Minitaur - an SNN accelerator designed for image inference and robotics applications [^4] There was also the n-Minitaur system(Kiselev et al., 2016) which was heavily based on Minitaur and was an adaptation of that system to support processing of the data from up to three spike-based sensors.. The architecture comprised 32 PEs with bespoke state and weight information caches. The neuron model was LIF with a decay rate based on the distance in time between the consecutive events. The system used broadcasting, instead of point-to-point connections. The authors chose a 16b fixed-point number representation for most of the parameters and values. The system implemented on Xilinx Spartan-6 device achieved 92% accuracy in MNIST benchmark, using an FF-FC network with 1010 neurons and was clocked at 75 MHz. Cheung et al. (2016) (Cheung et al., 2016) presented NeuroFlow - a complete framework for implementing dataflow-based NMAs [^5] This idea was first presented in (Cheung et al., 2012).. The design flow assumed that the necessary data was stored in the external memory, and processed by two sets of PEs: one for fetching the information and one for accumulating the spiking events and weights. The generated architectures supported a variety of neuron models (LIF, AdEx IF, IZH and HH), integration methods (Euler, RK4), synaptic input current kernels (exponential, alpha or custom) and a pair-based nearest neighbor STDP. The benchmarks included a simulation of 589824 neurons with varying numbers of synapses on four Xilinx Virtex-6 devices and a simple test of the STDP mechanics between two AdEx neurons. Wang et al. (2018) (Wang et al., 2018) presented an NMA for simulating mini- and hypercolumn models of the human neocortex. The system followed a hierarchical structure, where a network is implemented as a set of interconnected hypercolumns consisting of several minicolumns (up to 128), where every minicolumn comprised up to eight different types of heterogeneous neurons (100 per minicolumn). The neurons followed a two-compartment COBA LIF model, integrated with a 1 ms timestep. The representation of the weights and postsynaptic potentials was fixed-point 4b numbers, and the system supported axonal delay. The authors tested the system with a simulation of an auditory cortex divided into 100 channels sensitive to different sound frequencies. They mapped every channel to 100 hypercolumns, resulting in 10 thousand hypercolumns with 100 million LIF neurons. The authors claimed the maximum number of neurons and synapses supported by the architecture on an Altera Stratix-V to be 20 million and four trillion, respectively.
@@
Modern implementations. Han et al. (2020) (Han et al., 2020) presented an NMA with a single processing pipeline for LIF neuron dynamics. The authors introduced zero-biased ReLU-based offline training, as well as a hybrid updating algorithm, which, according to them, merged the event-driven and timed operation. They achieved it with separate event queues for a number of consecutive timesteps. The authors claimed the system supported 16384 neurons on Xilinx Kintex-7, and they tested it with the MNIST benchmark and FF-FC network (2058 neurons), for which it achieved an accuracy of 97.06% and was clocked at 200 MHz, with processing speed of 161 frames per second. Li et al. (2021) (Li et al., 2021) presented an NMA with a clock/event-driven hybrid update scheme - however, according to the Features described in Section 2.3, the system is a member of Class 3. It comprised 12 PEs with five computational cores (CC) in every PE, supporting CUBA LIF neurons with STDP-based learning capability. The four out of five CCs mentioned above could be shared with neighboring PEs. The architecture was implemented on Xilinx Virtex-7 and tested with thr MNIST benchmark and FF-FC (310 neurons) network topology with 16b floating-point representation, achieving 85.28% and the speed of 61FPS, clocked at 100 MHz. Liu et al. (2022) (Liu et al., 2022) presented FPGA-NHAP, designed with a simulation of FF topologies in mind. From an architectural point of view, the system was a single neural core that comprised 16 PEs and supported LIF and IZH neurons via two different pipelines, which were used to process neurons in a time-multiplexed manner. The network size was bound by the 14-bit addressing to 16384 neurons and 16.77 million synapses (with an A2A topology). The connectivity was stored in memory in a form of crossbars, and its size was fixed and architecturally bound. The authors chose a 16b signed fixed-point number representation. The system achieved the accuracy of 97.7% for LIF and 97.81% for IZH for the MNIST benchmark and 85.14% for LIF and 83.16% for IZH for the Fashion MNIST benchmark. Chen et al. (2024) (Chen et al., 2024b) presented SiBrain - a spatio-temporal parallel NMA targeting SCNNs. The architecture comprised 64 PEs. It achieved 99.08% accuracy for the MNIST benchmark with processing speed of 512 frames per second and 90.25% accuracy for the CIFAR-10 benchmark at 53 frames per second. The system was clocked at 200 MHz and implemented on Xilinx Virtex-7 2000T.
3.5. Class 4 (Traits: fully parallel, collocated computation and memory)
Figure 7. Simplified diagram of a Class 4 NMA.
@@ 3.5. Class 4 (Traits: fully parallel, collocated computation and memory)
Class 4 architectures support two Traits: fully parallel operation and collocation of memory and computation. They have the potential of updating the entire network the fastest out of all of the Classes, however due to the entire system - including the event processing circuitry - being contained on-chip, they are the most dependent on the amount of available resources on the FPGA platform. Early implementations. Roggen et al. (2003) (Roggen et al., 2003) presented an NMA for autonomous agents with on-the-fly reconfiguration capability. The authors chose a LIF neuron model, and the topology was a 2D array of 8x8 neurons. Architecture-wise, the system comprised 64 PEs with parameters kept on-chip, and it was clocked at 33 MHz on Altera APEX20K200E. Bellis et al. (2004) (Bellis et al., 2004) presented an NMA, which was a hardware port of a software-based simulator. The network comprised four LIF neurons and four synapses, and its main task was to solve the wall-following task for a an autonomous agent. Architecture-wise, we deduced that there were four PEs and that the system was clocked at 118.189 MHz on Xilinx Spartan-2, with a fixed-point representation of 5b width for weights and 7b for membrane potential. Upegui et al. (2005) (Upegui et al., 2005) presented a framework for design space exploration (DSE) for SNNs, based on the genetic algorithm approach (the authors called it the evolutionary algorithm). The authors used the Dynamic Partial Reconfiguration (DPR) for NMA reconfiguration. The architecture comprised a varying number of of PEs and a weight adjustment units and two fixed modules - encoder and decoder for input and output. The PEs and weight adjustment units were DPR-compliant, which could consist of single neurons, layers, or entire networks. The genetic algorithm was used to find the optimal network topology. The system supported a LIF neuron model and the STDP weight-adjusting method. Guerrero-Rivera et al. (2006) (Guerrero-Rivera et al., 2006) presented implementations of various logic designs for LIF neuronal dynamics with STDP and axonal delays using the Exact Integration method [^6] The authors thoroughly described it in another work (Guerrero-Rivera and Pearce, 2007).. The authors focused on deploying parallel, low-complexity neuron model implementations without resource sharing, which operated in a single-cycle manner. The architecture supported exponential decay for the LIF neurons, and alpha and beta synaptic kernels with STDP for synapses. The authors used the system to describe two classes of neurons in the mammalian olfactory bulb and deployed it on Xilinx Virtex-II Pro, clocked at 33 MHz. Shayani et al. (2008) (Shayani et al., 2008) presented the implementation of an SNN with Piecewise-Linear Approximation of Quadratic Integrate and Fire (PLAQIF) neurons. Their main goal was heterogeneous recurrent SNNs implementation, with support for a certain degree of structural plasticity. Architecture-wise, the system comprised 161 PEs with several synapse units per PE arranged in a column and interconnected in a daisy-chain fashion. The upstream packets (from soma) were unchanged and passed-through the synapses in the column, while downstream packets were processed by the synapse. The authors tested the implementation with a recurrent network of 161 neurons on the Xilinx Virtex-5 device clocked at 160 MHz.
@@
Early modern implementations. Johnston et al. (2010) (Johnston et al., 2010) presented an evolution-based SNN on FPGAs for robotics, designed with a HW/SW co-design methodology. They implemented the evolutionary part of the SNN a genetic algorithm, and added the STDP-based online learning for fine tuning the network obtained through genetic algorithm. The system functioned as an intelligent controller for a robotic system to solve autonomous navigation and obstacle avoidance. Architecture-wise, it relied on the network unit and embedded microcontroller to perform the update through the genetic algorithm. The authors tested the system on a Khepera robot(Mondada et al., 1999) with 10 LIF neurons and 16 synapses. The implementation was clocked at 100 MHz on a Xilinx Virtex-II Pro and used 16b floating-point arithmetic. Caron et al. (2011) (Caron et al., 2011) presented an implementation of the Hierarchical SNN (HSNN) based on the Oscillatory Dynamic Link Matcher (ODLM) algorithm, i.e., using LIF neurons to approximate the behavior of relaxation oscillators. They claimed that just like ODLM, the HSNN could perform image segmentation, matching, and monophonic sound source separation. Architecture-wise, the system comprised an HSNN unit, controller and communication units. The HSNN used a bit slice architecture composed of several columns connected in a ring buffer. Said slices were built off of a weight memory containing all of the presynaptic weights to the neuron instantiated by the column, and the synapse model unit, which added the synaptic weights to the neurons’ potential during spike propagation. Rossello et al. (2012) (RossellÓ et al., 2012) presented an NMA for stochastic SNNs. The architecture used a non-standard IF neuron model - the authors used a shift register where the single bit was being shifted depending on the signal from the synapses. The incoming spikes were only passed to the shifting logic if matched with the partial probabilities, realizing the stochastic operation. When the bit reached the MSB position, the spike was propagated further. The authors tested the system with a high-speed signal filtering and solving complex systems of linear equations tasks on Altera Cyclone-III, with a single layer of 100 output neurons. Iakymchuk et al. (2012) (Iakymchuk et al., 2012) proposed an NMA that efficiently implemented large networks on low-end hardware. The architecture supported a one millisecond biological timestep and featured a set of PEs that serially processed the input. The authors chose the SRM model for synapses and neurons and divided the system into three primary levels: neuron, layer, and network, which were realized by units consisting of a set of subunits from lower levels of the hierarchy. They utilized spike encoding with a specified frequency of spike trains, and chose 16b fixed-point arithmetic, with weights being drawn from a set of 30 16b fixed-point values. They tested the system on a Xilinx Spartan-3, with a pattern classification on 3x3-bit arrays using FF-FC network (9x9x4). Deng et al. (2014) (Deng et al., 2014) presented an NMA for a specific topology of an FF-FC network to achieve synchronization in neuron spiking between the layers, starting from uncorrelated firing rates. The architecture assumed using bespoke PEs per synapse and neuron. The authors chose the LIF neuron model and alpha kernel for the synaptic dynamics. The system achieved the intended synchronized firing between the layers on the Altera Stratix III clocked at 50 MHz, with an 8x8x8 network, achieving 50000x real-time operation. Wang et al. (2015) (Wang et al., 2015a) presented a general-purpose Liquid State Machine (LSM) network with a reconfigurable reservoir of randomly connected LIF neurons and pre-trained, task-specific fully-connected output layers. The authors implemented the reconfigurability of the reservoir’s neurons via power gating (turning off the unused parts of the FPGA). The plasticity was implemented with probabilistic weight updates and an induced teaching signal. The architecture relied on a set of PEs realizing the liquid neurons making up a reconfigurable unit and a set of PEs making up a trainable unit for the output functionality. The implementation presented in the paper used 135 liquid neurons (which can be sized down to 90), and 26 output neurons on a Xilinx Virtex-6. The number of output neurons depended on the output of the most demanding task the system was designed to perform. The accuracy for the MNIST benchmark was 98.6% for the largest available network (135 reservoir neurons and 26 output neurons). Farsa et al. (2015) (Farsa et al., 2015) presented an NMA for a single layer SNN performing function approximation. In a sense, it was a generalized classification task, as function computation was replaced by classifying input into various coded outputs. The authors chose the LIF neuron model for the system. The learning part was performed with MATLAB and SGD, which resulted in a set of weights that could be mapped to a single-layer WTA network. From the architectural point of view, the entire system comprised several PEs responsible for their bespoke neurons. The system was used to approximate the Gaussian force field as a test case and implemented on a Xilinx Zynq-7000. Gomar et al. (2016) (Gomar et al., 2016) presented an implementation of a neuron-synapse-neuron connection with additional astrocyte block connected to both neurons. Their main goal was to implement an astrocyte along the neurons efficiently. The authors employed an AdEx neuron model, the Kopel model for synapse and the Postnov model for astrocyte dynamics, resulting in six ODEs. As for the models, the authors optimized them by approximating the costly functions in the ODEs - exp and tanh - with Base-2 term and sign function, respectively. With the latter, the authors showed that Base-2 could also be used, but the sign was significantly cheaper implementation-wise. The implementation was tested on Xilinx Virtex-II. Lammie et al. (2018) (Lammie et al., 2018) did not focus on a specific architecture but rather on presenting an optimization approach for a specific NMA. The authors presented a way of utilizing a single vector of IZH neurons, arranged in a WTA topology with lateral inhibition, to recognize patterns after training them with STDP. The design was rate-based, and the spikes were Poissonian. The IZH neurons used bit shifts and adders to reduce the computational cost of multipliers. The STDP mechanism was simplified, where instead of holding the approximate values, it was approximated directly in circuitry via combination of OR gates. Heidarpur et al. (2019) (Heidarpur et al., 2019) presented a coordinate rotation digital computer (CORDIC)-based SNN based on IZH neurons. Despite the fact that the paper focused on the optimizations applied to the neuron computing pipeline rather than on the actual architecture choices, the authors suggested how to use this concept paired with STDP, which showed potential on how to create a larger NMA. The implemented network comprised 21 neurons, but the authors claimed they could potentially fit 110 neurons on a Xilinx Spartan-6 clocked at 183.4 MHz. Kuang et al. (2019) (Kuang et al., 2019) presented an NMA for realizing a specific SNN. The topology comprised three layers of neurons equal to the resolution of the input image. The operation assumed the output layer being made out of excitatory neurons, and the output is the firing pattern of this layer. The neurons were LIF with COBA synapses with exponential leakage and support for triplet-based STDP. The system was pipelined to increase the throughput. The authors tested the implementation with a simplified MNIST benchmark (5x5 instead of 28x28 pixels) for a network of 75 neurons that achieved 93% accuracy.

### v3

@@ 3.4 Class 3 (Traits: asynchronous network update)
Early implementations. Hellmich et al. (2004,2005) [128, 129] presented SNN Emulation Engine (SEE), which they claimed it could support up to 50 million IF neurons and 800 million synapses. The architecture organization relied on three FPGAs responsible for: (i)network configuration, monitoring and administration, (ii) computation of the target neurons addresses and (iii) updating neuron potentials and weight updates. A set of three PEs was implemented on the third device connected to three bespoke SDRAMs. The neuron update used a Bulirsch-Stoer integration method.
Figure 6: Simplified diagram of a Class 3 NMA.
@@
Ros et al. (2006) [130] presented RT-Spike, which was a SW/HW hybrid architecture for realizing SNNs with SRM-model-based neurons. The authors chose a 14b fixed-point number representation. They tested the system with a network of 1024 neurons on four PEs and achieved real-time operation on Xilinx Virtex-2000E at 25 MHz, with 4.16 million updates of the network state per second. Glackin et al. (2009) [131] presented Reconfigurable Architecture, which aimed at modeling the biological functions of the brain. The architecture comprised a set of PEs simultaneously simulating up to 250 neuron with COBA LIF and SRM neuron models support. Reconfiguring the network was performed on-the-fly, by changing the values in the BRAM memory via a custom DMA core. The implemented system performed edge detection on 512x512 px images on a Xilinx Virtex-4 device with four PEs, using an SCNN with 1.05 million neurons and 52.4 million synapses, reaching 1.05 seconds of inference time per image.
@@
Early modern implementations. Yang et al. (2011) [132] presented an NMA for explaining the computational capability of the early stage of the primate visual system. The input for the topology was a gray image transformed into a temporal firing rate in a given period. The architecture used four FPGAs, one per every image orientation feature, with 64 TDM PEs per device, allowing for one million LIF neurons and two million synapses. It was implemented on the Xilinx Virtex-4 device and clocked at 200 MHz. Neil et al. (2014) [133] presented Minitaur - an SNN accelerator designed for image inference and robotics applications [^4] There was also the n-Minitaur system[134] which was heavily based on Minitaur and was an adaptation of that system to support processing of the data from up to three spike-based sensors.. The architecture comprised 32 PEs with bespoke state and weight information caches. The neuron model was LIF with a decay rate based on the distance in time between the consecutive events. The system used broadcasting, instead of point-to-point connections. The authors chose a 16b fixed-point number representation for most of the parameters and values. The system implemented on Xilinx Spartan-6 device achieved 92% accuracy in MNIST benchmark, using an FF-FC network with 1010 neurons and was clocked at 75 MHz. Cheung et al. (2016) [80] presented NeuroFlow - a complete framework for implementing dataflow-based NMAs [^5] This idea was first presented in [81].. The design flow assumed that the necessary data was stored in the external memory, and processed by two sets of PEs: one for fetching the information and one for accumulating the spiking events and weights. The generated architectures supported a variety of neuron models (LIF, AdEx IF, IZH and HH), integration methods (Euler, RK4), synaptic input current kernels (exponential, alpha or custom) and a pair-based nearest neighbor STDP. The benchmarks included a simulation of 589824 neurons with varying numbers of synapses on four Xilinx Virtex-6 devices and a simple test of the STDP mechanics between two AdEx neurons. Wang et al. (2018) [135] presented an NMA for simulating mini- and hypercolumn models of the human neocortex. The system followed a hierarchical structure, where a network is implemented as a set of interconnected hypercolumns consisting of several minicolumns (up to 128), where every minicolumn comprised up to eight different types of heterogeneous neurons (100 per minicolumn). The neurons followed a two-compartment COBA LIF model, integrated with a 1 ms timestep. The representation of the weights and postsynaptic potentials was fixed-point 4b numbers, and the system supported axonal delay. The authors tested the system with a simulation of an auditory cortex divided into 100 channels sensitive to different sound frequencies. They mapped every channel to 100 hypercolumns, resulting in 10 thousand hypercolumns with 100 million LIF neurons. The authors claimed the maximum number of neurons and synapses supported by the architecture on an Altera Stratix-V to be 20 million and four trillion, respectively.
@@
Modern implementations. Han et al. (2020) [136] presented an NMA with a single processing pipeline for LIF neuron dynamics. The authors introduced zero-biased ReLU-based offline training, as well as a hybrid updating algorithm, which, according to them, merged the event-driven and timed operation. They achieved it with separate event queues for a number of consecutive timesteps. The authors claimed the system supported 16384 neurons on Xilinx Kintex-7, and they tested it with the MNIST benchmark and FF-FC network (2058 neurons), for which it achieved an accuracy of 97.06% and was clocked at 200 MHz, with processing speed of 161 frames per second. Li et al. (2021) [137] presented an NMA with a clock/event-driven hybrid update scheme - however, according to the Features described in Section 2.3, the system is a member of Class 3. It comprised 12 PEs with five computational cores (CC) in every PE, supporting CUBA LIF neurons with STDP-based learning capability. The four out of five CCs mentioned above could be shared with neighboring PEs. The architecture was implemented on Xilinx Virtex-7 and tested with thr MNIST benchmark and FF-FC (310 neurons) network topology with 16b floating-point representation, achieving 85.28% and the speed of 61FPS, clocked at 100 MHz. Liu et al. (2022) [138] presented FPGA-NHAP, designed with a simulation of FF topologies in mind. From an architectural point of view, the system was a single neural core that comprised 16 PEs and supported LIF and IZH neurons via two different pipelines, which were used to process neurons in a time-multiplexed manner. The network size was bound by the 14-bit addressing to 16384 neurons and 16.77 million synapses (with an A2A topology). The connectivity was stored in memory in a form of crossbars, and its size was fixed and architecturally bound. The authors chose a 16b signed fixed-point number representation. The system achieved the accuracy of 97.7% for LIF and 97.81% for IZH for the MNIST benchmark and 85.14% for LIF and 83.16% for IZH for the Fashion MNIST benchmark. Chen et al. (2024) [139] presented SiBrain - a spatio-temporal parallel NMA targeting SCNNs. The architecture comprised 64 PEs. It achieved 99.08% accuracy for the MNIST benchmark with processing speed of 512 frames per second and 90.25% accuracy for the CIFAR-10 benchmark at 53 frames per second. The system was clocked at 200 MHz and implemented on Xilinx Virtex-7 2000T.
3.5 Class 4 (Traits: fully parallel, collocated computation and memory)
Figure 7: Simplified diagram of a Class 4 NMA.
@@ 3.5 Class 4 (Traits: fully parallel, collocated computation and memory)
Class 4 architectures support two Traits: fully parallel operation and collocation of memory and computation. They have the potential of updating the entire network the fastest out of all of the Classes, however due to the entire system - including the event processing circuitry - being contained on-chip, they are the most dependent on the amount of available resources on the FPGA platform. Early implementations. Roggen et al. (2003) [140] presented an NMA for autonomous agents with on-the-fly reconfiguration capability. The authors chose a LIF neuron model, and the topology was a 2D array of 8x8 neurons. Architecture-wise, the system comprised 64 PEs with parameters kept on-chip, and it was clocked at 33 MHz on Altera APEX20K200E. Bellis et al. (2004) [141] presented an NMA, which was a hardware port of a software-based simulator. The network comprised four LIF neurons and four synapses, and its main task was to solve the wall-following task for a an autonomous agent. Architecture-wise, we deduced that there were four PEs and that the system was clocked at 118.189 MHz on Xilinx Spartan-2, with a fixed-point representation of 5b width for weights and 7b for membrane potential. Upegui et al. (2005) [142] presented a framework for design space exploration (DSE) for SNNs, based on the genetic algorithm approach (the authors called it the evolutionary algorithm). The authors used the Dynamic Partial Reconfiguration (DPR) for NMA reconfiguration. The architecture comprised a varying number of of PEs and a weight adjustment units and two fixed modules - encoder and decoder for input and output. The PEs and weight adjustment units were DPR-compliant, which could consist of single neurons, layers, or entire networks. The genetic algorithm was used to find the optimal network topology. The system supported a LIF neuron model and the STDP weight-adjusting method. Guerrero-Rivera et al. (2006) [143] presented implementations of various logic designs for LIF neuronal dynamics with STDP and axonal delays using the Exact Integration method [^6] The authors thoroughly described it in another work [144].. The authors focused on deploying parallel, low-complexity neuron model implementations without resource sharing, which operated in a single-cycle manner. The architecture supported exponential decay for the LIF neurons, and alpha and beta synaptic kernels with STDP for synapses. The authors used the system to describe two classes of neurons in the mammalian olfactory bulb and deployed it on Xilinx Virtex-II Pro, clocked at 33 MHz. Shayani et al. (2008) [145] presented the implementation of an SNN with Piecewise-Linear Approximation of Quadratic Integrate and Fire (PLAQIF) neurons. Their main goal was heterogeneous recurrent SNNs implementation, with support for a certain degree of structural plasticity. Architecture-wise, the system comprised 161 PEs with several synapse units per PE arranged in a column and interconnected in a daisy-chain fashion. The upstream packets (from soma) were unchanged and passed-through the synapses in the column, while downstream packets were processed by the synapse. The authors tested the implementation with a recurrent network of 161 neurons on the Xilinx Virtex-5 device clocked at 160 MHz.
@@
Early modern implementations. Johnston et al. (2010) [146] presented an evolution-based SNN on FPGAs for robotics, designed with a HW/SW co-design methodology. They implemented the evolutionary part of the SNN a genetic algorithm, and added the STDP-based online learning for fine tuning the network obtained through genetic algorithm. The system functioned as an intelligent controller for a robotic system to solve autonomous navigation and obstacle avoidance. Architecture-wise, it relied on the network unit and embedded microcontroller to perform the update through the genetic algorithm. The authors tested the system on a Khepera robot[147] with 10 LIF neurons and 16 synapses. The implementation was clocked at 100 MHz on a Xilinx Virtex-II Pro and used 16b floating-point arithmetic. Caron et al. (2011) [148] presented an implementation of the Hierarchical SNN (HSNN) based on the Oscillatory Dynamic Link Matcher (ODLM) algorithm, i.e., using LIF neurons to approximate the behavior of relaxation oscillators. They claimed that just like ODLM, the HSNN could perform image segmentation, matching, and monophonic sound source separation. Architecture-wise, the system comprised an HSNN unit, controller and communication units. The HSNN used a bit slice architecture composed of several columns connected in a ring buffer. Said slices were built off of a weight memory containing all of the presynaptic weights to the neuron instantiated by the column, and the synapse model unit, which added the synaptic weights to the neurons’ potential during spike propagation. Rossello et al. (2012) [149] presented an NMA for stochastic SNNs. The architecture used a non-standard IF neuron model - the authors used a shift register where the single bit was being shifted depending on the signal from the synapses. The incoming spikes were only passed to the shifting logic if matched with the partial probabilities, realizing the stochastic operation. When the bit reached the MSB position, the spike was propagated further. The authors tested the system with a high-speed signal filtering and solving complex systems of linear equations tasks on Altera Cyclone-III, with a single layer of 100 output neurons. Iakymchuk et al. (2012) [150] proposed an NMA that efficiently implemented large networks on low-end hardware. The architecture supported a one millisecond biological timestep and featured a set of PEs that serially processed the input. The authors chose the SRM model for synapses and neurons and divided the system into three primary levels: neuron, layer, and network, which were realized by units consisting of a set of subunits from lower levels of the hierarchy. They utilized spike encoding with a specified frequency of spike trains, and chose 16b fixed-point arithmetic, with weights being drawn from a set of 30 16b fixed-point values. They tested the system on a Xilinx Spartan-3, with a pattern classification on 3x3-bit arrays using FF-FC network (9x9x4). Deng et al. (2014) [151] presented an NMA for a specific topology of an FF-FC network to achieve synchronization in neuron spiking between the layers, starting from uncorrelated firing rates. The architecture assumed using bespoke PEs per synapse and neuron. The authors chose the LIF neuron model and alpha kernel for the synaptic dynamics. The system achieved the intended synchronized firing between the layers on the Altera Stratix III clocked at 50 MHz, with an 8x8x8 network, achieving 50000x real-time operation. Wang et al. (2015) [152] presented a general-purpose Liquid State Machine (LSM) network with a reconfigurable reservoir of randomly connected LIF neurons and pre-trained, task-specific fully-connected output layers. The authors implemented the reconfigurability of the reservoir’s neurons via power gating (turning off the unused parts of the FPGA). The plasticity was implemented with probabilistic weight updates and an induced teaching signal. The architecture relied on a set of PEs realizing the liquid neurons making up a reconfigurable unit and a set of PEs making up a trainable unit for the output functionality. The implementation presented in the paper used 135 liquid neurons (which can be sized down to 90), and 26 output neurons on a Xilinx Virtex-6. The number of output neurons depended on the output of the most demanding task the system was designed to perform. The accuracy for the MNIST benchmark was 98.6% for the largest available network (135 reservoir neurons and 26 output neurons). Farsa et al. (2015) [153] presented an NMA for a single layer SNN performing function approximation. In a sense, it was a generalized classification task, as function computation was replaced by classifying input into various coded outputs. The authors chose the LIF neuron model for the system. The learning part was performed with MATLAB and SGD, which resulted in a set of weights that could be mapped to a single-layer WTA network. From the architectural point of view, the entire system comprised several PEs responsible for their bespoke neurons. The system was used to approximate the Gaussian force field as a test case and implemented on a Xilinx Zynq-7000. Gomar et al. (2016) [154] presented an implementation of a neuron-synapse-neuron connection with additional astrocyte block connected to both neurons. Their main goal was to implement an astrocyte along the neurons efficiently. The authors employed an AdEx neuron model, the Kopel model for synapse and the Postnov model for astrocyte dynamics, resulting in six ODEs. As for the models, the authors optimized them by approximating the costly functions in the ODEs - exp and tanh - with Base-2 term and sign function, respectively. With the latter, the authors showed that Base-2 could also be used, but the sign was significantly cheaper implementation-wise. The implementation was tested on Xilinx Virtex-II. Lammie et al. (2018) [155] did not focus on a specific architecture but rather on presenting an optimization approach for a specific NMA. The authors presented a way of utilizing a single vector of IZH neurons, arranged in a WTA topology with lateral inhibition, to recognize patterns after training them with STDP. The design was rate-based, and the spikes were Poissonian. The IZH neurons used bit shifts and adders to reduce the computational cost of multipliers. The STDP mechanism was simplified, where instead of holding the approximate values, it was approximated directly in circuitry via combination of OR gates. Heidarpur et al. (2019) [156] presented a coordinate rotation digital computer (CORDIC)-based SNN based on IZH neurons. Despite the fact that the paper focused on the optimizations applied to the neuron computing pipeline rather than on the actual architecture choices, the authors suggested how to use this concept paired with STDP, which showed potential on how to create a larger NMA. The implemented network comprised 21 neurons, but the authors claimed they could potentially fit 110 neurons on a Xilinx Spartan-6 clocked at 183.4 MHz. Kuang et al. (2019) [157] presented an NMA for realizing a specific SNN. The topology comprised three layers of neurons equal to the resolution of the input image. The operation assumed the output layer being made out of excitatory neurons, and the output is the firing pattern of this layer. The neurons were LIF with COBA synapses with exponential leakage and support for triplet-based STDP. The system was pipelined to increase the throughput. The authors tested the implementation with a simplified MNIST benchmark (5x5 instead of 28x28 pixels) for a network of 75 neurons that achieved 93% accuracy.